Two largest stadiums with the largest capacities are Marshfield stadium and Bowden stadium. Marshfield stadium has a capacity of 6,053 more than Bowden. If the combined capacity for the two stadiums is 210,355, find the capacity for each stadium.
Answer by elima(1433)

Two largest stadiums with the largest capacities are Marshfield stadium and Bowden stadium. Marshfield stadium has a capacity of 6,053 more than Bowden. If the combined capacity for the two stadiums is 210,355, find the capacity for each stadium.
B + M =210355
M = B + 6053
B + (B + 6053)=210355
2B+6053=210355
2B=204302
B=102151
M=102151+6053=108204
:)